In nuclear fission, a nucleus splits roughly in half.

A. what is the potential 4.00x10-14m from a fragment that has 50 protons in it? (p=+1.602x10-19C)

B. what is the potential energy of a similarly charged fragment at this distance?


1
Expert's answer
2021-04-23T11:05:45-0400

A. Charge of the fragment:

"Q = 50 \;protons = 50 \times 1.6 \times 10^{-19} \;C"

Distance from the fragment:

"r = 4.00 \times 10^{-14} \;m"

Electric force constant:

"k = 9 \times 10^9 \;Nm^2/C^2"

The electric potential of a point charge:

"V = \frac{kQ}{r} \\

V = \frac{9 \times 10^9 \times 50 \times 1.6 \times 10^{-19}}{4.00 \times 10^{-14}} \\

= 1.80 \times 10^6 \;V"

B. The relationship between potential and electrical potential energy is:

"V = \frac{U}{q} \\

U = qV \\

U = 50 \times 1.80 \times 10^6 \\

= 9.0 \times 10^7 \;eV \\

= 90 \;MeV"